OpenStudy (kj4uts):

There are values of t so that sin t = .35 and cos t = .6 True or false. I don't understand what this is asking?

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sin^2(t) + \cos^2(t) = 1\]

jimthompson5910 (jim_thompson5910):

this is the pythagorean identity

OpenStudy (kj4uts):

It would be false sin^2t+cos^2t not equal 1?

jimthompson5910 (jim_thompson5910):

what did you get instead of 1?

OpenStudy (kj4uts):

0.7987... I think that's right?

jimthompson5910 (jim_thompson5910):

you need to calculate 0.35^2+0.6^2

jimthompson5910 (jim_thompson5910):

Notice I just replaced sine and cosine with their respective values

OpenStudy (kj4uts):

0.35^2+0.6^2=0.4825

jimthompson5910 (jim_thompson5910):

that is not equal to 1, so the overall statement is false

jimthompson5910 (jim_thompson5910):

to make it true, you would have to change sin(t) to 0.8 sin(t) = 0.8 cos(t) = 0.6 makes sin^2(t) + cos^2(t) = 1 true. This isn't the only way to fix it of course.

OpenStudy (kj4uts):

I understand now so if it equals 1 (is this for any problem like this) its true and if not its false like this question?

jimthompson5910 (jim_thompson5910):

for this type of problem, yes, and always think back to the pythagorean identity the goal is to make it true

OpenStudy (kj4uts):

So false here. Thanks for your help again your awesome!

jimthompson5910 (jim_thompson5910):

yes false

OpenStudy (kj4uts):

I wish I had a math teacher like you :)

jimthompson5910 (jim_thompson5910):

I'm sure your math teacher is fine. Perhaps he's just too overwhelmed? Not sure.